OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

\[\sin ^ {2}x or sinx\]

OpenStudy (anonymous):

could this be \(7\sin^2(x)-14\sin(x)+2=-7\)?

OpenStudy (anonymous):

yea, sorry

OpenStudy (anonymous):

sinx=1

OpenStudy (anonymous):

start with \[7\sin^2(x)-14\sin(x)+7=0\] then divide by 7 and get \[\sin^2(x)-2\sin(x)=1\] a perfect square this gives \[(\sin(x)-1)^2-0\] and so \[\sin(x)=1\]

OpenStudy (anonymous):

Oh! Seems simple enough:) thanks

OpenStudy (anonymous):

i made a typo, should have written \[\sin^2(x)-2\sin(x)-1=0\]